OpenStudy (yueyue):

Solve each of the following. Please give both the general solution and the specific solutions in the given interval. 1. 2sin(2theta)+√3=0 interval [0,2π)

OpenStudy (anonymous):

\[2\sin2 \theta+\sqrt{3}=0\]

OpenStudy (anonymous):

\[\sin2\theta=-\sqrt{3/4}\]

OpenStudy (anonymous):

\[2\sin(2\theta)+√3=0 \rightarrow 2\sin(2\theta)=-√3 \rightarrow \sin(2\theta)=-√3/2\ \[\sin(2\theta)=-√3/2 \rightarrow \sin(2\theta)=- \sin \frac{\pi}{3} =\sin (\pi+\frac{\pi}{3})\] \[ \sin(2\theta)= =\sin (\frac{4\pi}{3}) \rightarrow 2\theta =\frac{4\pi}{3} \rightarrow \theta = \frac{4\pi}{3 \times 2}\] \[\rightarrow \theta = \frac{2\pi}{3 }\]

OpenStudy (anonymous):

@yueyue

OpenStudy (yueyue):

@dpasingh Alright thank you but do you have any idea what it means when it mentions the general solution and the specific solutions?